# ViTextVQA: A Large-Scale Visual Question Answering Dataset for Evaluating Vietnamese Text Comprehension in Images

### v2

@@ 3.2.2 Answer Analysis
POS Tags in Answer
Similar to the question, POS tags in the answer not only help to better understand the grammatical structure but also provide important information about the content and meaning of the answer in the VQA task.
Figure 10: Distribution of Postag in answer.
Figure 11: Distribution of NER in answer.
In Figure 11, we can notice the richness and diversity of the Vietnamese language used in the answer. Noun (N) make up the largest proportion of the data with approximately 70,000 instances, these tags often appear to designate objects, people, or objects associated with the image. This shows a focus on providing information about what answers the image-related question.
In addition, the Verb (V) also plays an important role. With nearly 30,000 and ranked after noun (N), it shows that annotators are also interested in describing the actions or states of objects in the image, making the answers more vivid.
Punctuation and symbols (CH) and Numeral/Quantity (M) appear the third most with about 25,000 times. This is understandable because the dataset focuses on text in the image, and dots often appear in phone numbers (e.g., 0909.888.888), building addresses with punctuation, or on advertising banners. In contrast, answers about the price of product, number of houses, etc. often contain a numeral or quantity factor (M).
Adjectives (A) are often used to express the characteristics or conditions of the object or entity described in the answer. With around 9,000 occurrences, they show that, even if they are just tokens from the OCR that the annotator sees in the image, they still contain specific messages and linguistic colors to convey meaning best.
Moreover, several other tags, while they may be in the minority, significantly contribute to enriching the diversity of answers in the dataset.
NER Tags in Answer
NER tags also play an important role in identifying specific entities mentioned in the answer, help us better understand the variety of entities emphasized in answering image-based questions, and provide information that the model can pay attention to.
Figure 11 provides insight into how language is used in answers annotated by native Vietnamese speakers, with particular attention to entities such as places, organizations, and individuals. The majority of words in the answer are not classified as concrete entities (O) similar to the question. However, we also see the emergence of specific entities such as places (B-LOC, I-LOC), organizations (B-ORG, I-ORG), and individuals (B-PER), although not too prominent with less than 4,000 or even less. This diversifies the content of the answer, helping the model understand and generalize about possible real-life human cases.
3.2.3 Object Analysis
Object in Image
@@ 4 Methodology
The common text-based VQA approach employs three key components to tackle image-based questions: Visual Features Extraction, Textual Features Extraction, and OCR System (see Figure 14). In this framework, the Visual Features Extraction extracts vital information from images, while the Textual Features Extraction encapsulates questions into features vectors. Meanwhile, OCR System undertakes the crucial task of recognizing and extracting text portrayed in images. Through the seamless integration of these three components, the text-based VQA method can comprehend and answer questions from images with enhanced accuracy and efficiency.
@@ 4.2 Visual Features Extraction
ViT introduced by Dosovitskiy et al. [41], a pre-trained image model based on transformer architecture, has been trained on a large amount of image data, playing an important role in the English baseline models mentioned in Section 4.4. In the baseline adaptation to Vietnamese, we kept the same visual feature extraction using ViT. Meanwhile, When adapting Vietnamese to conduct experiments on the ViTextVQA dataset with baseline models based on CNN-LM based method, we chose VinVL [60], a model based on Faster R-CNN [40] to extract object features. This choice also brings convenience when we use VinVL to analyze object in Section 3.2.3.
@@ 4.4 Baseline Models
There are numerous effective methods have emerged globally for addressing VQA task. To evaluate the difficulty of the ViTextVQA dataset, we choose M4C [63], LaTr [8], PreSTU [10], BLIP-2 [43] and SaL [64] according to the historical development (Figure 15) of VQA methods and use them as baselines to perform experiments. Note that these baselines were originally designed to support English, but we adapted them to Vietnamese while still retaining their essence.
@@
M4C VinVL mBERT SwinTextSpotter CNN-LM
@@
For a comprehensive overview of the baseline models, refer to Table 5 along with additional details provided in Sections 4.1 and 4.2 regarding their pre-trained model. Furthermore, Section 4.3 details the OCR system, and Section 2.2 elaborates on the different types of methods employed by these models.
M4C: The emergence of Facebook’s TextVQA dataset [4] catalyzed the development of a pioneering approach called Multimodal Multi-Copy Mesh (M4C) [63], marking the first instance of treating VQA as a generated answer. M4C employs BERT [32] for question embedding, utilizing its encoder module as a multimodal encoder to embed diverse features, including object and region features in images, question features from BERT, visual features of scene texts, and features of previous tokens in answers. In a departure from conventional classification approaches, M4C views VQA as a generative task and introduces the Dynamic Pointer Network, as proposed by Hu et al. [63]. This network seamlessly integrates features of embedded tokens in the vocabulary with those of scene texts. In particular, M4C makes decisions at each step of the answer generation process, determining whether to select tokens from the vocabulary or scene texts.
LaTr: The Layout-Aware Transformer (LaTr), as introduced by Biten et al. [8], adopts the encoder-decoder transformer architecture, specifically building upon T5 [65]. The model comprises three primary blocks. The initial module is dedicated to a Language Model trained for document layouts, focusing on text and layout details. The second module employs spatial embedding to embed tokens related to scene text along with their positional information, represented by OCR tokens bounding boxes. Utilizing the Vision Transformer (ViT) [41], the third module extracts visual features. Subsequently, these embedding features from the three modalities serve as input for the pre-trained transformer. The encoder then learns a representation that harmonizes information from these modalities. This learned representation is subsequently employed by a decoder for analysis and generation of an output, typically in the form of an answer.
PreSTU: Pre-Training for Scene-Text Understanding [10] stands out from other models in tackling a common task by incorporating a unique approach. To ensure that models grasp spatial information in scene text and standardize the target output sequence during training, the authors arranged OCR text by their positions, sequentially from the top-left to the bottom-right. They then concatenated the sorted texts using a T5 separator ¡/s¿. Afterward, they randomly split the OCR text into two segments, utilizing the first part as additional input and the second part as the target. Next, the research team applied the T5 architecture to conduct pre-training on a large amount of scene text image data. The goal is to help the model, which is a language model, be able to learn information related to scene text in the image. During the pretraining and finetuning phase, ViT [41] is used to extract image features, assisting PreSTU in predicting the target answer based on the question and OCR text contained in the image.
BLIP-2: BLIP-2 [43] short for “Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models”, is a model that can understand and generate text related to images and do so in languages. It combines a vision transformer, a query transformer, and a large language model (LLM). In the structure of the model, they frozen image encoders and large language models (LLMs), meaning these components do not learn during training, query transformer (Q-Former) that acts as a lightweight bridge between the image encoder and LLM. This model was trained using the weights of ViT and T5 on multiple VQA datasets.
SaL: Most current Text-based VQA methods focus on improving network design or pre-training tasks. However, these approaches often treat text in images as a simple sequence, neglecting the lack of semantic connections between many words and the complex spatial relationships between them. Fang et al. [64] proposed a novel method called “Separate and Locate” (SaL) that addresses these issues. SaL explores two special modules, including “Text Semantic Separation” (TSS) and “Spatial Circle Position” (SCP). TSS module identifies words that have true semantic relationships and separates them from those that are simply positioned close together. This helps the model focus on meaningful connections between words. SCP module introduced a new way to represent the spatial relationships between words. Instead of relying on simple 1D positioning, SCP uses a “spatial circle” embedding that captures the relative distances and directions between words, providing a more accurate understanding of their layout. By combining these two modules, SaL achieves significant improvements in accuracy on TextVQA [4] and ST-VQA [5] datasets, outperforming state-of-the-art methods without requiring any pre-training itself.
@@ 4.5.1 Exact Match
EM=Count of questions predicted exactlyCount of $\textbf{EM}=\frac{\text{Count of questions predicted exactly}}{\text{Count of questions}}$ (1)
@@ 4.5.2 F1-score
5 Experiment and Results
5.1 Experimental Configuration
The baseline models were trained using the Adam optimization algorithm [66] on an RTX 4090 GPU with 24GB of memory. Each model underwent 10 epochs, with approximately 7 hours required for completion using ViT-LM based method and 6 hours using CNN-LM based method. Hyperparameters for the baseline model were set as follows: a learning rate of 3.0e-5, dropout rate of 0.2, batch size of 16, and weight decay of 1.0e-4.
@@ 5.2 Main Results
Table 6 presents a comprehensive of the results achieved by individual models, showcasing their F1-Scores and EM metrics. The findings underscore significant disparities in performance among the various models, providing valuable insights into their respective strengths and weaknesses. Notably, within the baseline models, PreSTU large version based on ViT-LM method stands out as the top performer, showcasing an impressive F1-Score of 44.93 and an EM of 22.64. Moreover, the table illuminates the substantial performance enhancements attainable in VQA task through the utilization of large-scale pre-trained models.
@@ 6 Results Analysis
6.1 ViT5 and Effect of OCR Text as Question Suffix
During the experiment, we discovered that models using ViT5 as their backbone achieve better performance when the given question is suffixed with OCR text from images compared to the scenario in which only the question is passed to the VQA models as in previous studies. In this respect, we are interested in further exploring the performance that models such as BLIP-2, LaTR, PreSTU, and SaL perform when using OCR text as an important part of the “context” for the question. This experiment was conducted by adding and removing OCR text along with the question, thereby helping us with more detailed assessment of the capabilities of the models.
Figure 16: Performance of models include OCR text and without OCR text scenario.
Figure 16 (left) illustrates the F1-score of the models when using and not using OCR text. Models trained with question extended by OCR text consistently outperformed models that did not use OCR text. In particular, PreSTU model (large) is a standout, with an F1 score increasing from 16.95 (without OCR) to 44.93 (with OCR), showing a significant increase in performance when integrating OCR text in the training process. This positive trend is also repeated in other models.
Similarly, with the EM score which is shown in Figure 16 (right), it can be seen that models trained with OCR text consistently demonstrated almost twice the accuracy compared to models that did not use OCR. This indicates that integrating OCR text into the training process not only improves performance but also enhances the accuracy of the models.
Overall, the use of OCR text to expand the question as a “context” for models using ViT5 as their backbone has proven effective and has the potential to improve the performance of these models on VQA task. In other words, as answers in the ViTextVQA dataset are text or scene text from images, providing OCR text along with questions makes the VQA models in our experiments find it easier to “select” appropriate tokens from OCR text for determining the answers.
@@ 6.2 Effect of OCR System Performance and Challenge
Figure 18: Results of models based on OCR system performance.
@@ 6.5 Effect of OCR Arrangement
Table 8: Results of models with different OCR arrangements. $\triangle$ indicates the increase (↑) or decrease (↓) and compares with the original (*).
Model Not Sort Confidence Score Sort Top-Left To Bottom-Right Sort F1-Score (%) EM (%) F1-Score (%) EM (%) F1-Score (%) EM (%) M4C $\triangle$ 25.46* - 10.26* - 18.41 $\downarrow 7.05$ 8.29 $\downarrow 1.97$ 30.04 $\uparrow 4.58$ 11.60 $\uparrow 1.34$ BLIP-2(base) $\triangle$ 36.87* - 13.59* - 35.21 $\downarrow 1.66$ 12.97 $\downarrow 0.62$ 37.78 $\uparrow 0.91$ 15.01 $\uparrow 1.42$ BLIP-2(large) $\triangle$ 38.47* - 15.58* - 36.74 $\downarrow 1.73$ 14.15 $\downarrow 1.43$ 42.16 $\uparrow 3.69$ 18.56 $\uparrow 2.98$ Latr(base) $\triangle$ 40.41* - 17.85* - 38.76 $\downarrow 1.65$ 16.32 $\downarrow 1.53$ 43.13 $\uparrow 2.72$ 20.42 $\uparrow 2.57$ Latr(large) $\triangle$ 42.34* - 20.94* - 39.72 $\downarrow 2.62$ 18.88 $\downarrow 2.06$ 44.07 $\uparrow 1.73$ 22.27 $\uparrow 1.33$ SaL(base) $\triangle$ 43.49* - 20.11* - 43.66 $\uparrow 0.17$ 20.28 $\uparrow 0.17$ 44.89 $\uparrow 1.40$ 20.97 $\uparrow 0.86$ SaL(large) $\triangle$ 43.39* - 21.48* - 43.71 $\uparrow 0.32$ 20.47 $\downarrow 1.01$ 44.74 $\uparrow 1.35$ 21.18 $\downarrow 0.30$ PreSTU(base) $\triangle$ 41.74 $\downarrow 2.07$ 19.15 $\downarrow 1.70$ 40.09 $\downarrow 3.72$ 17.74 $\downarrow 3.11$ 43.81* - 20.85* - PreSTU(large) $\triangle$ 42.85 $\downarrow 2.08$ 21.40 $\downarrow 1.24$ 40.13 $\downarrow 4.80$ 18.43 $\downarrow 4.21$ 44.93* - 22.64* - Avg 39.46 17.81 37.38 16.39 41.73 19.28
@@
SaL model: This model exhibits a relatively stable performance regardless of whether it is sorted by top-left to bottom-right or confidence score or not sort. This phenomenon can be attributed to the distinctive characteristics of its two specialized modules, TSS and SCP. These modules provide the model with the capability to pay attention to tokens that have actual significance within OCR text. Consequently, it becomes evident that the SaL model is not heavily reliant on the specific location of tokens in OCR text. This characteristic reduces the dependence of the model on the order of OCR, contributing to its consistent performance across different arrangement types.

### v3

@@ 3.3.2 Answer Length
3.3.3 Object in Image
@@ 3.3.4 Object in Question
3.4 Comparison with Other Visual Question Answering Datasets
Figure 10: Comparison of question length among VQA datasets.
In order to contextualize our dataset within the broader landscape of Visual Question Answering (VQA), we compare it with several widely used VQA datasets based on key textual characteristics: question length and answer length. These attributes influence model performance, as shorter questions may be simpler to parse, while longer questions can provide more context or require complex reasoning. Similarly, answer length can affect model expressiveness and comprehension.
ViTextVQA is distinguished by its ability to accommodate substantially longer questions compared to existing datasets (see Figure 10), making it particularly suitable for tasks that require deeper reasoning and detailed query formulation. In contrast to VQAv2, TextVQA, and ViVQA, which predominantly feature shorter questions, ViTextVQA maintains a consistent distribution of questions exceeding 15 words. This characteristic enhances its applicability to complex visual question-answering (VQA) tasks that demand more comprehensive language understanding. While OpenViVQA provides a broader range of question lengths, ViTextVQA further extends this diversity by sustaining a high frequency of longer queries, thereby facilitating more intricate reasoning processes.
Figure 11: Comparison of answer length among VQA datasets.
In terms of answer length, ViTextVQA allows for substantially longer responses compared to VQAv2, ViVQA, and TextVQA, providing greater flexibility for VQA tasks (see Figure 11). While OpenViVQA maintains a more balanced answer length distribution, ViTextVQA surpasses it by supporting answers of up to 50 words, enabling more detailed and comprehensive textual responses. This capability reinforces ViTextVQA as a valuable resource for advancing research in VQA, particularly for tasks that require complex reasoning based on textual information extracted from images.
@@ 4 Methodology
Our text-based VQA method employs three key components to tackle image-based questions: Visual Features Extraction, Textual Features Extraction, and OCR System (see Figure 12). In this method, the Visual Features Extraction extracts vital information from images, while the Textual Features Extraction encapsulates questions into features vectors. Meanwhile, OCR System undertakes the crucial task of recognizing and extracting text portrayed in images. Through the seamless integration of these three components, the text-based VQA method can comprehend and answer questions from images with enhanced accuracy and efficiency.
We propose ViTextBLIP-2 is an extension of BLIP-2 (see Figure 13), designed to overcome computational and data limitations in Vietnamese OCR-based Visual Question Answering (OCR-VQA). Instead of training from scratch, the model focuses on the second stage of BLIP-2, optimizing language generation by leveraging a frozen Large Language Model (LLM). To enhance Vietnamese text comprehension, ViTextBLIP-2 integrates additional modules, including an OCR system and an image captioning model. During training, only the Q-Former is optimized, while other components remain frozen to reduce computational costs.
Figure 13: The structure of ViTextBLIP-2.
Image Encoder: A pretrained Vision Transformer (ViT) is employed and kept frozen to retain its learned visual representations.
$E_{\text{img}}=\text{ViT}(I)$ (1)
where $E_{\text{img}}$ is the image embedding, and $I$ is the input image.
OCR System: The model incorporates SwinTextSpotter, maintaining its original architecture and parameters. The extracted text tokens and embeddings are directly passed to Q-Former.
$\{T_{\text{OCR}},E_{\text{OCR}}\}=\text{SwinTextSpotter}(I)$ (2)
where $T_{\text{OCR}}$ represents OCR tokens and $E_{\text{OCR}}$ denotes OCR embeddings.
Captioning Model: The Qwen2-VL model is utilized to generate automatic image captions, providing additional contextual information. This module remains frozen during training.
$C=\text{Qwen2VL}(I)$ (3)
where $C$ represents the generated caption.
Q-Former: This component is trained to integrate multimodal features from the Image Encoder, OCR System, and Captioning Model using Self-Attention and Cross-Attention mechanisms.
$Q=\text{Q-Former}(E_{\text{img}},E_{\text{OCR}},Q_{\text{init}})$ (4)
Large Language Model (LLM): The ViT5 Phan et al. [58] model, a pretrained Vietnamese LLM, is incorporated as the language generation component. Its parameters are kept frozen, while the system optimizes its capability to generate responses based on the processed multimodal embeddings.
$A=\text{ViT5}(Q,Q_{\text{ques}},T_{\text{OCR}},C)$ (5)
where $A$ is the generated answer, and $Q_{\text{ques}}$ is the input question.
@@ 4.2 Visual Features Extraction
ViT introduced by Dosovitskiy et al. [39], a pre-trained image model based on transformer architecture, has been trained on a large amount of image data, playing an important role in the English baseline models mentioned in Section 5.1. In the baseline adaptation to Vietnamese, we kept the same visual feature extraction using ViT. Meanwhile, When adapting Vietnamese to conduct experiments on the ViTextVQA dataset with baseline models based on CNN-LM based method, we chose VinVL [56], a model based on Faster R-CNN [38] to extract object features. This choice also brings convenience when we use VinVL to analyze object in Section LABEL:obj_analysis.
@@ 5.1 Baseline Models
M4C VinVL mBERT-cased SwinTextSpotter CNN-LM
@@
To evaluate the difficulty of the ViTextVQA dataset, we selected several state-of-the-art VQA models as baselines according to their historical development 14. These include M4C [60], which pioneered treating VQA as a generative task using BERT for question embedding and introducing the Dynamic Pointer Network for integrating vocabulary and scene text features; LaTr [8], which adopts an encoder-decoder transformer architecture based on T5 with three primary modules focusing on document layouts, spatial embedding, and visual features extraction; PreSTU [10], which uniquely arranges OCR text by position and uses T5 architecture for pre-training on scene text image data; BLIP-2 [41], which combines frozen image encoders and large language models with a query transformer bridge; and SaL [61], which addresses the neglect of semantic connections and spatial relationships between words through its Text Semantic Separation and Spatial Circle Position modules. While these baselines were originally designed for English, we adapted them to Vietnamese while preserving their core functionalities. Table 5 provides a comprehensive overview of these models’ feature extraction methods and OCR systems, with additional details on their pre-trained models and methodologies available in the referenced sections.
5.2 Large Language Models (LLMs)
Currently, large language models (LLMs) are entering a race to develop advanced techniques to solve the Visual Question Answering (VQA) problem. Leading models such as GPT-4o [2], Gemini-1.5-Flash [3], and Qwen2-VL [62]are continuously being improved to enhance performance in analyzing, understanding, and reasoning from image data, helping to improve user experience. These advancements not only help improve visual information processing capabilities but also create many potential applications in academia, industry, and daily life.
GPT-4o enhances VQA capabilities through image analysis, excelling in object recognition, OCR, and chart interpretation. Gemini-1.5 Flash, Google’s multimodal chatbot, processes images and text simultaneously, delivering real-time answers to visual queries with optimized speed. Qwen2-VL, an open-source multimodal model from Alibaba Research, handles complex tasks like chart analysis and multilingual text recognition, utilizing Naive Dynamic Resolution and Multimodal Rotary Position Embedding (M-ROPE) for improved visual-textual integration.
5.3 Experimental Configuration
The baseline models were trained using the Adam optimization algorithm [63] on an RTX 4090 GPU with 24GB of memory. Each model underwent 10 epochs, with approximately 7 hours required for completion using ViT-LM based method and 6 hours using CNN-LM based method. Hyperparameters for the baseline model were set as follows: a learning rate of $3.0e^{-5}$, dropout rate of 0.2, batch size of 16, and weight decay of $1.0e^{-4}$.
@@ 5.4.1 Exact Match
EM=Count of answers predicted exactlyCount of $\textbf{EM}=\frac{\text{Count of answers predicted exactly}}{\text{Count of answers}}$ (6)
@@ 5.5 Main Results
ViTextBLIP-2 (Ours) base ViT-LM 53.95 25.48
Table 6 presents a comprehensive overview of the results achieved by individual models, showcasing their F1-score and EM metrics. The findings underscore significant disparities in performance among the various models, providing valuable insights into their respective strengths and weaknesses. Notably, within the baseline models, the large version of PreSTU, based on the ViT-LM method, stands out as the top performer with an impressive F1-score of 44.93 and an EM of 22.64. More notably, the proposed ViTextBLIP-2 model achieves an outstanding F1-score of 53.95 and an EM of 25.48, demonstrating a remarkable improvement over the baseline models. These results highlight the superior potential of the proposed model in enhancing the performance of text-based VQA tasks in Vietnamese.
5.6 Performance of LMMs Compared to Human
Table 7: LLM results.
Model F1-score (%) EM (%)
0-shot 1-shot 3-shot 5-shot 0-shot 1-shot 3-shot 5-shot
GPT-4o 55.51 69.55 76.67 75.62 25.00 45.00 56.00 55.00
Gemini-1.5-flash 57.75 59.09 62.37 64.40 26.00 29.00 35.00 39.00
QwenVL-7b 53.80 54.10 54.85 54.74 28.00 36.00 36.00 35.00
ViTextBLIP-2 56.43 26.00
Human 96.42 88.00
We randomly selected 100 samples from the test set for evaluation. Specifically, the group used large language models (LLMs) to perform zero-shot and few-shot (including 1, 3, and 5 shots) measurements and compared the results with human-generated answers for the same 100 samples.
From Table 7, it can be seen that the most advanced large language models (LLMs) today, such as GPT-4o, Gemini-1.5-flash, and QwenVL-7b, still lag behind humans in terms of providing answers in visual question answering (VQA) tasks based on images containing text, especially in Vietnamese. Although there is an improvement from 0-shot to few-shot, the F1-score and Exact Match (EM) scores of these models are still much lower than humans. Humans achieved scores of 96.42 (F1) and 88.00 (EM), while the models only reached scores from 53.80 to 75.62 (F1) and 25.00 to 55.00 (EM). This shows that LLMs are still not capable of generating answers as accurate and nuanced as humans in VQA tasks involving Vietnamese text.
5.7 Compare with Other Dataset
Table 8: Performance comparison of different models.
Model Version Method Type CIDEr Avg BLEU
MCAN base ViT-LM 1.0613 0.1699
LoRRA base CNN-RNN 0.8005 0.1349
M4C base CNN-LM 1.5073 0.2941
FST base CNN-LM 0.6141 0.1050
QuMLPAG base CNN-LM 1.7082 0.2651
MLPAG base CNN-LM 1.6104 0.2739
ViTextBLIP-2 (our) base ViT-LM 3.2129 0.4717
Table 8 presents the results of traditional methods such as MCAN, LoRRA, and M4C based on standard evaluation metrics like CIDEr and Avg BLEU on the OpenViVQA dataset. M4C, with a CIDEr score of 1.5073 and Avg BLEU of 0.2941, pioneered integrating reading comprehension with visual question answering (VQA). The proposed models (FST, QuMLPAG, and MLPAG) were also evaluated, where QuMLPAG achieved a strong performance with a CIDEr score of 1.7082 and average BLEU of 0.2651.
Notably, ViTextBLIP-2 outperformed all models with a CIDEr score of 3.2129 and Avg BLEU of 0.4717, surpassing both traditional VQA models and the proposed approaches. This result highlights its superiority not only on the authors’s dataset but also on other benchmarks.
@@ 6.1 Effect of OCR System Performance and Challenge
Figure 16: Results of models based on OCR system performance.
@@ 6.4 Effect of OCR Arrangement
In other words, since we often observe that scene text tends to follow a Top-Left To Bottom-Right pattern, we decided to explore whether sorting the OCR outputs in different ways might improve the model’s performance. By experimenting with various sorting strategies, we wanted to see if organizing the text in a specific order could help the VQA model process the information more effectively.
@@
Table 10: Results of models with different OCR arrangements. $\triangle$ indicates the increase (↑) or decrease (↓) compared with the original (*).
Model Not Sorted Confidence Score Sorted Top-Left To Bottom-Right Sorted
F1-score (%) EM (%) F1-score (%) EM (%) F1-score (%) EM (%)
M4C 25.46* 10.26* 18.41 (↓7.05) 8.29 (↓1.97) 30.04 (↑4.58) 11.60 (↑1.34)
BLIP-2 base 36.87* 13.59* 35.21 (↓1.66) 12.97 (↓0.62) 37.78 (↑0.91) 15.01 (↑1.42)
BLIP-2 large 38.47* 15.58* 36.74 (↓1.73) 14.15 (↓1.43) 42.16 (↑3.69) 18.56 (↑2.98)
LaTr base 40.41* 17.85* 38.76 (↓1.65) 16.32 (↓1.53) 43.13 (↑2.72) 20.42 (↑2.57)
LaTr large 42.34* 20.94* 39.72 (↓2.62) 18.88 (↓2.06) 44.07 (↑1.73) 22.27 (↑1.33)
SaL base 43.49* 20.11* 43.66 (↑0.17) 20.28 (↑0.17) 44.89 (↑1.40) 20.97 (↑0.86)
SaL large 43.39* 21.48* 43.71 (↑0.32) 20.47 (↓1.01) 44.74 (↑1.35) 21.18 (↓0.30)
PreSTU base 41.74 (↓2.07) 19.15 (↓1.70) 40.09 (↓3.72) 17.74 (↓3.11) 43.81* 20.85*
PreSTU large 42.85 (↓2.08) 21.40 (↓1.24) 40.13 (↓4.80) 18.43 (↓4.21) 44.93* 22.64*
ViTextBLIP-2 (Ours) 51.16 (↓1.35) 24.10 (↓0.73) 53.95 (↑1.44) 25.48 (↑0.65) 52.51* 24.83*
Avg 40.62 18.45 39.04 17.30 42.81 19.83
@@
SaL and ViTextBLIP-2: These two models exhibit stable performance regardless of the OCR token sorting method, whether arranged top-left to bottom-right, by confidence score, or unsorted. For the SaL model, this stability is due to its TSS (Token Selection Strategy) and SCP (Semantic Context Processing) modules, which help it focus on meaningful tokens rather than their spatial arrangement. Similarly, the ViTextBLIP-2 model achieves comparable robustness through its Q-Former module, which collaborates with the OCR system to filter out noise and extract essential features. By leveraging these mechanisms, both models minimize dependence on OCR token order, ensuring consistent performance across different sorting strategies.
@@
6.5 Ablation Study
Table 11: Ablation study on different model components.
Image Encoder OCR System Captioning Model F1-score (%) EM (%)
0.5251* 0.2483*
0.3914 (↓0.1337) 0.1694 (↓0.0789)
0.3953 (↓0.1298) 0.1794 (↓0.0689)
0.1417 (↓0.3834) 0.0469 (↓0.2014)
0.4167 (↓0.1084) 0.1882 (↓0.0601)
0.3668 (↓0.1583) 0.1332 (↓0.1511)
0.3512 (↓0.1739) 0.1267 (↓0.1216)
0.0016 (↓0.5235) 0.0000 (↓0.2483)
The goal of this experiment is to explore the performance of the ViTextBLIP-2 model when combining different components, including the Image Encoder, OCR System, and Image Captioning Model. Understanding the impact of each component on model performance helps optimize image and text processing systems to improve recognition and caption generation accuracy. Details of this ablation analysis are presented in Table 11.
Full Component Integration (✓ ✓ ✓): When all three components are included in the model, the highest performance is achieved, with an F1-score of 0.5251 and an EM score of 0.2483. This indicates that the full integration of all systems enables the model to maximize information utilization, leading to the best performance. Combining all three components allows the model to not only understand the image but also generate more accurate answers.
Image Encoder and OCR System Only (✓ ✓ ✗): When the Image Encoder is combined with the OCR System but without the Image Captioning Model, the F1-score drops to 0.3914 and the EM score decreases to 0.1694. Although OCR assists in extracting textual information, the absence of an image captioning model limits the model’s ability to generate semantically rich descriptions, resulting in lower performance.
Image Encoder and Image Captioning Model Only (✓ ✗ ✓): When the Image Encoder is combined with the Image Captioning Model but without OCR, the performance further declines, with an F1-score of 0.3953 and an EM score of 0.1794. While the model can generate descriptive captions for images, the lack of OCR support prevents it from recognizing text within the image, leading to missing textual information that could enhance the response. The absence of OCR reduces the model’s ability to handle text-heavy images.
Image Encoder Only (✓ ✗ ✗): When only the Image Encoder is used without other components, the performance drops significantly, with an F1-score of just 0.1417 and an EM score of 0.0469. This suggests that the Image Encoder alone is insufficient for generating accurate answers, as it lacks both OCR-derived text information and image-captioning semantic understanding.
No Components Used (✗ ✗ ✗): When none of the modules are utilized, the F1-score and EM score drop to their lowest values, at just 0.0016 and 0.0000, respectively. This demonstrates that the model is completely unable to produce meaningful results. The reason is that without any image-related information, the model cannot process the context needed to understand the query or generate accurate predictions. This clearly highlights the importance of image-based information for the ViTextBLIP-2 model to function effectively.